# Alternative Approaches for Computing Highest-Density Regions

### v1

@@ Appendix B Supporting Details for Hyperparameter Choices
Figure 23: Performance results in all evaluated scenarios for $n=100$ and for different values of the hyperparameter $\epsilon$ in $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(1.74-0.26\log(n)\right)=1.72$.
@@
Figure 24: Performance results in all evaluated scenarios for $n=500$ and for different values of the hyperparameter $\epsilon$ in $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(1.74-0.26\log(n)\right)=1.13$.
@@
Figure 25: Performance results in all evaluated scenarios for $n=1000$ and for different values of the hyperparameter $\epsilon$ in $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(1.74-0.26\log(n)\right)=0.95$.
@@
Figure 26: Performance results in the Dirichlet scenario for all sample sizes and for different values of the hyperparameter $\epsilon$ in $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF. The optimal choice reflects the following heuristic: $\epsilon=\exp\left(-1.22-0.23\log(n)\right)$.
@@
Figure 27: Performance results in all evaluated scenarios for $n=50$ and for different values of the hyperparameter $\epsilon$ in $M^{\text{PCop}}_{3}$:$\epsilon$-CDF. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(1.60-0.41\log(n)\right)\approx 1$.
@@
Figure 28: Performance results in all evaluated scenarios for $n=100$ and for different values of the hyperparameter $\epsilon$ in $M^{\text{PCop}}_{3}$:$\epsilon$-CDF. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(1.60-0.41\log(n)\right)=0.75$.
@@
Figure 29: Performance results in all evaluated scenarios for $n=500$ and for different values of the hyperparameter $\epsilon$ in $M^{\text{PCop}}_{3}$:$\epsilon$-CDF. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(1.60-0.41\log(n)\right)=0.39$.
@@
Figure 30: Performance results in all evaluated scenarios for $n=1000$ and for different values of the hyperparameter $\epsilon$ in $M^{\text{PCop}}_{3}$:$\epsilon$-CDF. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(1.60-0.41\log(n)\right)=0.29$.
@@
Figure 31: Performance results in the Dirichlet scenario for all sample sizes and for different values of the hyperparameter $\epsilon$ in $M^{\text{PCop}}_{3}$:$\epsilon$-CDF. Gray vertical line reflects the optimal choice given by $\epsilon=0.02$.
@@ Appendix C Simulation results for all scenarios
In this section, we report the full set of simulation results (for all scenarios and sample sizes) that were omitted in the main paper for the sake of space.

### v2

@@ Appendix B Supporting Details for Hyperparameter Choices
Figure 22: Performance results of $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF in all evaluated scenarios for $n=100$ and for different values of the hyperparameter $\epsilon$, with $\epsilon\in\{0.1,0.5,\dots,9.5,10.0\}$. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(1.74-0.26\log{n}\right)=1.72$.
@@
Figure 23: Performance results of $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF in all evaluated scenarios for $n=500$ and for different values of the hyperparameter $\epsilon$, with $\epsilon\in\{0.1,0.5,\dots,9.5,10.0\}$. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(1.74-0.26\log{n}\right)=1.13$.
@@
Figure 24: Performance results of $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF in all evaluated scenarios for $n=1000$ and for different values of the hyperparameter $\epsilon$, with $\epsilon\in\{0.1,0.5,\dots,9.5,10.0\}$. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(1.74-0.26\log{n}\right)=0.95$.
@@
Figure 25: Performance results of $M^{\text{NPCop}}_{3}$:$\epsilon$-CDF in the Dirichlet scenario for all sample sizes and for different values of the hyperparameter $\epsilon$, with $\epsilon\in\{0.001,0.010,0.020,\dots,0.290,0.300\}$. The optimal choice reflects the following heuristic: $\epsilon=\exp\left(-1.22-0.23\log{n}\right)$.
@@
Figure 26: Performance results of $M^{\text{PCop}}_{3}$:$\epsilon$-CDF in all evaluated scenarios for $n=50$ and for different values of the hyperparameter $\epsilon$, with $\epsilon\in\{0.01,0.02,0.05,0.10,0.20,0.30,\dots,4.0,5.0\}$. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(1.60-0.41\log{n}\right)\approx 1$.
@@
Figure 27: Performance results of $M^{\text{PCop}}_{3}$:$\epsilon$-CDF in all evaluated scenarios for $n=100$ and for different values of the hyperparameter $\epsilon$, with $\epsilon\in\{0.01,0.02,0.05,0.10,0.20,0.30,\dots,4.0,5.0\}$. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(1.60-0.41\log{n}\right)=0.75$.
@@
Figure 28: Performance results of $M^{\text{PCop}}_{3}$:$\epsilon$-CDF in all evaluated scenarios for $n=500$ and for different values of the hyperparameter $\epsilon$, with $\epsilon\in\{0.01,0.02,0.05,0.10,0.20,0.30,\dots,4.0,5.0\}$. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(1.60-0.41\log{n}\right)=0.39$.
@@
Figure 29: Performance results of $M^{\text{PCop}}_{3}$:$\epsilon$-CDF in all evaluated scenarios for $n=1000$ and for different values of the hyperparameter $\epsilon$, with $\epsilon\in\{0.01,0.02,0.05,0.10,0.20,0.30,\dots,4.0,5.0\}$. Gray vertical line reflects the following heuristic: $\epsilon=\exp\left(1.60-0.41\log{n}\right)=0.29$.
@@
Figure 30: Performance results of $M^{\text{PCop}}_{3}$:$\epsilon$-CDF in the Dirichlet scenario for all sample sizes and for different values of the hyperparameter $\epsilon$, with $\epsilon\in\{0.001,0.005,0.010,0.015,\dots,0.195,0.200\}$. Gray vertical line reflects the optimal choice given by $\epsilon=0.02$.
@@ Appendix C Simulation results for all scenarios
In this section, we report the full set of simulation results (for all scenarios and sample sizes) that were omitted in the main paper for the sake of space. All reported results are summarized in terms of mean (and standard deviation) across $1000$ independent Monte Carlo (MC) replicates.
@@
Figure 31: Global comparison between the evaluated measures in terms of the Matthews correlation coefficient (MCC) metric, for scenarios S1-S16 and sample sizes $n\in\{50,100,250,500,1000\}$.